OpenStudy (anonymous):

The limit lim (√(25+h)-5)/h h->0 represents the derivative of some function f(x) at some number a. Find f and a.

OpenStudy (anonymous):

This is the derivative. Solve it and integrate result.

OpenStudy (anonymous):

integrate the result*

OpenStudy (anonymous):

I should clarify. You'll have to just observe what the function is based on the definition of the derivative. Namely, \[f'(x)=\lim_{\Delta x \to 0}\frac{f(x+\Delta x)-f(x)}{\Delta x}.\]

OpenStudy (anonymous):

or if you know the definition of the derivative: f`(a)=lim (f(a+h)-f(a))/h here f(a+h)=sqrt(25+h) f(a)=5

OpenStudy (anonymous):

@Andras, you have to uniquely find \(f(x)\) and the value it is being evaluated, \(a.\) Also, please do not post full answers. It's outlined against in the Code of Conduct.

OpenStudy (anonymous):

a=0

OpenStudy (anonymous):

but f is my problem

OpenStudy (anonymous):

Why do you think it's \(a=0?\)

OpenStudy (anonymous):

i evaluated the limit

OpenStudy (anonymous):

Sorry havent used the site for ages, but I dont think I have given full answer.

OpenStudy (anonymous):

Maybe it would help to put this in better terms: \[f'(a)=\lim_{h \to 0}\frac{f(a+h)-f(a)}{h}.\] Can you figure out what \(a\) is now? @Andras, yes, but I thought that was your intention previously. Sorry if I was wrong in presuming such.

OpenStudy (anonymous):

a=5?

OpenStudy (anonymous):

That's close. Try again.

OpenStudy (anonymous):

√25

OpenStudy (anonymous):

Nope. \(a\) is a positive number. Try this, equate the two: \[\lim_{h \to 0}\frac{f(a+h)-f(a)}{h}=\lim_{h \to 0}\frac{\sqrt{25+h}-5}{h}.\] Can you see it now?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So, what must \(a\) be?

OpenStudy (anonymous):

square root of 25+h

OpenStudy (anonymous):

Not quite. That's \(f(a+h).\) If \(f(a+h)=\sqrt{25+h}\), what is \(a?\)

OpenStudy (anonymous):

square root of 25

OpenStudy (anonymous):

Why do you think it's \(\sqrt{25}?\)

OpenStudy (anonymous):

because i plugged in \[\sqrt{25+h}\] in f(a+h) ?

OpenStudy (anonymous):

You can't plug in anything into \(f(a+h);\) it is already defined for you in \(f(a+h)=\sqrt{25+h}\). You have to find the \(a\) such that it's a true statement.

OpenStudy (anonymous):

how can i find it?

OpenStudy (anonymous):

Well, \(f(a+h)=\sqrt{a+h}=\sqrt{25+h}.\) Do you see what \(a\) must be in order for the statement to be true?

OpenStudy (anonymous):

25?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Now, what is \(f(x)?\)

OpenStudy (anonymous):

can you give me a hint ?

OpenStudy (anonymous):

Well, we've seen that \(f(a+h)=\sqrt{a+h}.\) So, what would that look like if you let \(x=a+h?\)

OpenStudy (anonymous):

\[\sqrt{25+h}\]

OpenStudy (anonymous):

Not quite. We're letting \(x=a+h\). So what would \(f(x)\) look like, since we know \(f(a+h)=\sqrt{a+h}?\)

OpenStudy (anonymous):

\[\sqrt{25+x}\]

OpenStudy (anonymous):

Nope. You're getting close. Imagine we put \(x\) where \(a+h\) is. Do you know what \(f(x)\) is now?

OpenStudy (anonymous):

f (\[\sqrt{25+h}\])

OpenStudy (anonymous):

Well, that's not quite it. We're trying to look for the function which looks like \(f(a+h)=\sqrt{a+h}.\) What does \(f(x)\) have to be?

OpenStudy (anonymous):

can you give me examples ?

OpenStudy (anonymous):

Yes. If we have \(f(a+h)=a+h\), then \(f(x)=x.\) Or, if we have \(f(a+h)=(a+h)^2\), then \(f(x)=x^2.\) Another example: If \(f(a+h)=e^{a+h},\) then \(f(x)=e^x.\) Also, if \(f(a+h)=\sqrt[3]{a+h},\) then \(f(x)=\sqrt[3]{x}.\)

OpenStudy (anonymous):

√x?

OpenStudy (anonymous):

Correct! Great job.

OpenStudy (anonymous):

so f(x)=√x, and a=25?

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

thanks very very thanks

OpenStudy (anonymous):

You're welcome! :D